SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Oct 17, 2006)

Attendees: Debbie Bard, Rainer Bartoldus, Al Eisner, Selina Li, Dong Su, Mark Tibbetts.

Ops

  • Last Friday Su Dong tidied up the lab considerably, unplugging all cables, sorting them and plugging them back in. The ethernet cables still need to be organised. The fiber splitter is still in place, and this should be removed at some point. There should be some testing done once all this is finished, to check things are still working correctly.

EMT

  • Experts info has been updated with Mark's details.
  • Instructions for EMT prescaling are here.

L3

  • Rainer was paged this morning as the LST was taking some test data and had problems with 40% deadtime at a 4000Hz rate. The slow-down was caused by the continuous L3 damage printout from dataflow damage and IFR missing contributions. Rainer created a new trigger mask (see here) to enable LST to take test data without the deadtime problems.

AOB

  • An updated list of current Trigger tasks is being prepared, and can be found here. Everyone is free to add details or more things to do.
  • The next meeting will be in 2 weeks time, on Oct 31st.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.